31
May
2013
19:01

There is currently an issue with some BT connections. The problem is outside of our control hower as soon as we have an indication of a time to repair we will let you know. Sorry for any inconvenience this is causing.

Posted by Phil McGillivray

29
May
2013
17:04

I am pleased to say that the core router upgrades went to plan today – we believe no packets were injured during this period.

Posted by Kevin Hones

27
May
2013
20:29

The exploit that has been occurring on and off all weekend was diagnosed to be a series of malformed TCP port 80 packets – i.e. probably trying to exploit weaknesses in webservers. The FireBrick team have issued new firmware to protect against this and we have accordingly updated all or FireBricks except our core routers.

Customers using FB2500 or FB2700 models with default settings will see their bricks have updated firmware automatically. Anyone who has disabled auto-updates may have had ongoing problems all weekend – the solution is to update to the current factory firmware.

We are planning an update to our core network routers, which should not be disruptive but must be considered an ‘at-risk’ period. Until this has been completed, the network status will remain on Amber. Please see the planned maintenance page for details for the core upgrade.

Posted by Kevin Hones

25
May
2013
05:00

There currently seems to be an exploit attack which is causing FireBricks to spontaneously reboot. Fortunately they do so very quickly, but this has resulted in many brief outages as parts of our core network are resetting.
The exploit has been understood and the FireBrick team have already issued a new release of firmware to avoid it. We will be deploying this across our core as soon as feasible.

Posted by Kevin Hones

23
May
2013
11:13

Apologies to our customers for the brief connectivity glitch to all colo equipment in bluesquare 2. This was because I dislodged a cable to our firewall/edge router ‘grappa’ during installation of new machines. Sorry for this – I promise to be more careful.

Posted by Yolanda Zyczynska

07
May
2013
15:43

Update: This issue has now been resolved

We have been notified by our upstream suppliers that there is a technical fault within the 02 core network which is currently being investigated. Some Be line customers may be experiencing packet loss.

We update as soon as more information is available.

Posted by Yolanda Zyczynska

Current Status

    GoodNo current issues

Status Archive